# Scalable Composite Likelihood Estimation of Probit Models with Crossed Random Effects

### v2

@@ Abstract
Crossed random effects structures arise in many scientific contexts. They raise severe computational problems with likelihood computations scaling like $N^{3/2}$ or worse for $N$ data points. In this paper we develop a new composite likelihood approach for crossed random effects probit models. For data arranged in $R$ rows and $C$ columns, the likelihood function includes a very difficult $R+C$ dimensional integral. The composite likelihood we develop uses the marginal distribution of the response along with two hierarchical models. The cost is reduced to $\mathcal{O}(N)$ and it can be computed with $R+C$ one dimensional integrals. We find that the commonly used Laplace approximation has a cost that grows superlinearly. We get consistent estimates of the probit slope and variance components from our composite likelihood algorithm. We also show how to estimate the covariance of the estimated regression coefficients. The algorithm scales readily to a data set of five million observations from Stitch Fix with $R+C>700{,}000$.
Keywords: Composite Likelihood; High-Dimensional Binary Data; Probit Regression; Scalability.
@@ 1 Introduction
In this paper we develop a scalable approach to mixed effects modeling with a probit link and a crossed random effects error structure. The common notation for mixed effects models incorporates the fixed and random effects through a formula such as $X\bm{\beta}+Z\bm{b}$ using matrices $X$ and $Z$ of known predictors and unknown coefficient $\bm{\beta}$ and $\bm{b}$ where $\bm{b}$ happens to be random. This formulation is simple and elegant but it hides an important practical difference. When the random effects have a hierarchical structure then computations usually scale well to large problems. However, when the random effects have a crossed structure, then computations can easily grow superlinearly, which greatly restricts the use of these models. For instance, computing the maximum likelihood estimator in a Gaussian model with crossed random effects by the usual algorithms includes linear algebra with a cost that grows at least as fast as $N^{3/2}$ given $N$ data points (Gao and Owen, 2020), while remaining $\mathcal{O}(N)$ in the hierarchical setting. The binary regression setting we consider here has the further difficulty of a very high dimensional integral that is not present for a model with a Gaussian response. Maximizing the likelihood with the expectation-maximization algorithm raises the same intractability problems as direct likelihood maximization because the integral appearing in the expectation step has the same very high dimension as the integral appearing in the likelihood.
@@
In an era where the size $N$ of data sets is growing rapidly it is not possible to use algorithms with a cost of $\Omega(N^{3/2})$. Ideally the cost should be $\mathcal{O}(N)$. We are motivated by electronic commerce problems with large data sets. A company might have customers $i=1,\dots,R$, to which it sells items $j=1,\dots,C$. It would then be interested in modeling how a response $Y_{ij}$ depends on some predictors $\bm{x}_{ij}\in\mathbb{R}^{p}$. If they do not account for the fact that $Y_{ij}$ and $Y_{is}$ are correlated due to a common customer $i$ or that $Y_{ij}$ and $Y_{rj}$ are correlated due to a common item $j$, they will not get efficient estimates of the coefficients in their model. Much more seriously, they will not get reliable standard errors for those estimates. In a setting where accounting for random effects is computationally impossible we expect that many users will simply ignore them, getting very naive variance estimates and finding too many things significant.
@@
A typical feature of data in our motivating applications is a very sparse sampling. Only $N\ll RC$ of the possible $(\bm{x}_{ij},Y_{ij})$ values are observed. There is generally no simple structure in the pattern of which $(i,j)$ pairs are observed. While we use electronic commerce as our example, large sparsely sampled crossed data sets arise in other areas too.
@@
One of the main open questions in generalized linear mixed models, like the ones we consider here, is at what rate do the estimators of the model parameters converge? Intuitively one might expect $\mathcal{O}(N^{-1/2})$ or $\mathcal{O}(\min\{R,C\}^{-1/2})$. In some settings estimators for different parameters converge at different rates. See Jiang (2013) for some discussion. It was only comparatively recently that consistency of maximum likelihood estimation was proved in some generalized linear mixed models. This is also due to Jiang (2013).
In this paper we focus on binary responses, $Y_{ij}\in\{0,1\}$. In commerce applications, such responses may be more important than real-valued ones. Examples include whether customer $i$ purchased item $j$, or returned it or reported satisfaction with it. In addition to the numerical algebra that underlies the $\Omega(N^{3/2})$ cost of standard algorithms this setting brings a second difficulty of an $R+C$ dimensional integral over random effect values in the likelihood. It is very common to consider a logistic regression for such binary responses. In this paper we consider a probit regression. The probit model is especially suitable because it is typical to use Gaussian models for the random effects. The probit and logit link functions are nearly proportional outside of tail regions (Agresti, 2002, pp. 246–247), and so they often give similar results. We have some further remarks about the logistic case in our conclusions.
We consider two crossed random effects. There is a vector $\bm{a}\in\mathbb{R}^{R}$ with elements $a_{i}$ and another vector $\bm{b}\in\mathbb{R}^{C}$ with elements $b_{j}$. Conditionally on $\bm{a}$ and $\bm{b}$, the $Y_{ij}$ are independent with
$\Pr(Y_{ij}=1\!\mid\!\bm{a},\bm{b})=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}+a_{i}+b_{j}),$ (1)
where $\Phi(\cdot)$ is the standard normal cumulative distribution function. In these models, random effects are typically assumed to be uncorrelated, $\bm{a}\sim\mathcal{N}(0,\sigma^{2}_{A}I_{R})$ independently of $\bm{b}\sim\mathcal{N}(0,\sigma^{2}_{B}I_{C}),$ where $I_{n}$ is the $n\times n$ identity matrix, see, for example, McCullagh and Nelder (1989, page 444). The probit model has a representation in terms of latent variables $\varepsilon_{ij}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,1)$, via
$Y_{ij}=1\{\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}+a_{i}+b_{j}+\varepsilon_{ij}>0\},$ (2)
where $1\{E\}$ is the indicator function of the event $E$. The probability (1) and the likelihoods derived from it are all conditional on the values of $\bm{x}_{ij}$. By combining (1) and (2) we find that marginally
$\Pr(Y_{ij}=1)=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})$ (3)
for $\bm{\gamma}=\bm{\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}$. The marginal model (3) can be fit in $\mathcal{O}(N)$ computation, but we then need estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to get the scale right. While $\mathrm{sign}(\bm{\gamma}_{k})=\mathrm{sign}(\beta_{k})$ for $k=1,\dots,p$ confidence intervals for $\bm{\gamma}_{k}$ based on model (3) would be naive if they did not account for the dependence among the responses.
We can similarly find that
$\Pr(Y_{ij}=1\!\mid\!\bm{a})$ $=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{A}+u_{i}),\quad\text{and}$ (4)
$\Pr(Y_{ij}=1\!\mid\!\bm{b})$ $=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{B}+v_{j})$ (5)
where $\bm{u}\sim\mathcal{N}(0,\tau^{2}_{A}I_{R})$ and $\bm{v}\sim\mathcal{N}(0,\tau^{2}_{B}I_{C})$ for $\tau^{2}_{A}=\sigma^{2}_{A}/(1+\sigma^{2}_{B})$, $\tau^{2}_{B}=\sigma^{2}_{B}/(1+\sigma^{2}_{A})$, $\bm{\gamma}_{A}=\bm{\gamma}(1+\tau^{2}_{A})^{1/2}$ and $\bm{\gamma}_{B}=\bm{\gamma}(1+\tau^{2}_{B})^{1/2}$. We fit models (3), (4) and (5) together to estimate our probit model, using the latter two to estimate $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. Fitting models (4) and (5) involve simpler integrals than (1) since the latent variable representations of these models
$Y_{ij}=1\{\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{A}+u_{i}+\varepsilon_{ij}>0\}\quad\text{and}\quad Y_{ij}=1\{\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{B}+v_{j}+\varepsilon_{ij}>0\}$
have hierarchical (not crossed) error structures. This is where we are able to replace an $R+C$-dimensional integral by $R+C$ univariate ones.
We call our method ‘all-row-column’ (ARC). The name comes from model (3) that uses all the data at once, model (4) that combines likelihood contribution from within each row and model (5) that combines likelihood contributions from within each column. This approach is a new form of composite likelihood (Lindsay, 1988; Varin et al., 2011). An earlier version of composite likelihood that was applied to crossed random effects is discussed in Bellio and Varin (2005). They considered a standard pairwise likelihood which is however not scalable and thus inappropriate for our problem. Instead, our ARC method does not only consistently estimates $\bm{\beta}$, $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ but it is also scalable having $\mathcal{O}(N)$ cost.
An outline of this paper is as follows: Section 2 develops further notation and properties for the models we combine. Section 3 presents our ARC method and develops a robust sandwich estimate of the variance of the regression parameter estimates. Section 4 presents some simulations of the ARC method. We also compare it to an infeasible oracle method that estimates $\bm{\gamma}$ from (3) and then adjusts the estimate using the true values of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. A second comparison is to the most popular frequentist method for fitting generalized linear models with random effects, that is the first-order Laplace approximate maximum likelihood estimator described in Skaug (2002) and implemented in the popular R (R Core Team, 2024) package lme4 (Bates et al., 2015). Section 5 illustrates our proposed method on real electronic commerce data provided by Stitch Fix. Section 6 gives some bibliographic references focussing on the challenges that the crossed random effects setting brings. Section 7 has our conclusions. Finally, there is a supplement with proofs of consistent estimation, more references, consideration of pairwise composite likelihood and a more detailed discussion of the simulation results.
2 Probit Regression with Crossed Random Effects
As described in the introduction we have responses $Y_{ij}\in\{0,1\}$, covariates $\bm{x}_{ij}\in\mathbb{R}^{p}$ and random effects $\bm{a}\sim\mathcal{N}(0,\sigma^{2}_{A}I_{R})$, $\bm{b}\sim\mathcal{N}(0,\sigma^{2}_{B}I_{C})$. The random effects are independent of each other, and we condition on the observed values of the covariates $\bm{x}_{ij}$, so we then treat them as non-random. We write $\mathcal{S}\subset\{1,\dots,R\}\times\{1,\dots,C\}$ for the set of $(i,j)$ pairs where $(\bm{x}_{ij},Y_{ij})$ was observed. We also work conditionally on $\mathcal{S}$. In our motivating applications the pattern of observation/missingness could well be informative. Addressing that issue would necessarily require information from outside the data. Furthermore, the scaling problem is still the subject of current research in the non-informative missingness setting. Therefore we consider estimation strategies without taking account of missingness.
We choose to study the probit model (1) where, conditionally on $\bm{a}$ and $\bm{b}$, $Y_{ij}\in\{0,1\}$ are independent with $\Pr(Y_{ij}=1\!\mid\!\bm{a},\bm{b})=\Phi(\bm{x}_{ij}^{\mathsf{T}}\beta+a_{i}+b_{j})$. The full likelihood for $\bm{\theta}=(\bm{\beta}^{\mathsf{T}},\sigma^{2}_{A},\sigma^{2}_{B})^{\mathsf{T}}$ is the integral
$L(\bm{\theta})=\sigma_{A}^{-R}\sigma_{B}^{-C}\int_{\mathbb{R}^{R+C}}L(\bm{\beta}\!\mid\!\bm{a},\bm{b})\prod_{i=1}^{R}\varphi\Bigl{(}\frac{a_{i}}{\sigma_{A}}\Bigr{)}\prod_{j=1}^{C}\varphi\Bigl{(}\frac{b_{j}}{\sigma_{B}}\Bigr{)}\,\mathrm{d}\bm{a}\,\mathrm{d}\bm{b},$ (6)
where $\varphi(\cdot)$ is the standard normal probability density function, and $L(\bm{\beta}\!\mid\!\bm{a},\bm{b})$ is the conditional likelihood of $\bm{\beta}$ given the random effects. The conditional likelihood we need is
$L(\bm{\beta}\!\mid\!\bm{a},\bm{b})=\prod_{(i,j)\in\mathcal{S}}\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}+a_{i}+b_{j})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}-a_{i}-b_{j})^{1-y_{ij}}.$
An option to alleviate the computational cost of likelihood computations is provided by composite likelihood methods (Lindsay, 1988). Composite likelihoods are inference functions constructed by multiplying component likelihoods. See Varin et al. (2011) for a review. Several articles discussed composite likelihood estimation of models for dependent binary data; see, for example, Heagerty and Lele (1998), Cox and Reid (2004) or Kuk (2007). All these works consider a particular type of composite likelihood called pairwise likelihood that is built taking the product of bivariate marginal probabilities. Bellio and Varin (2005) investigated pairwise likelihood inference in logistic regression with crossed random effects. Their pairwise likelihood involves all the pairs of correlated observations, that is those pairs that share the row- or the column-random effect.
Pairwise likelihood is not attractive in this setting. The number of pairs of data points far exceeds the $\mathcal{O}(N)$ constraint. In a setting with $R$ rows, the average row has $N/R$ elements in it. If all the rows had that many elements then the number of pairs would be $\Omega(R(N/R)^{2})=\Omega(N^{2}/R)$. Unequal numbers of observations per row can only increase this count. As a result the cost must be $\Omega(\max(N^{2}/R,N^{2}/C))$. In our motivating settings, both $R$ and $C$ are $o(N)$ so the cost cannot then be $\mathcal{O}(N)$. A possible remedy to the lack of scalability would be to use only $\mathcal{O}(N)$ randomly sampled pairs of correlated observations. Although sampling might be a valid option, we decided to avoid this approach. First, we prefer estimates that do not depend on unnecessary randomness in the algorithm. Second, some preliminary simulations reported in Section S2 of the Supplementary Material indicate that the composite likelihood discussed in the next section satisfy the $\mathcal{O}(N)$ constraint with a statistical efficiency comparable to that of the pairwise likelihood without sampling of the pairs, which is however not feasible to compute in the high-dimensional applications that motivate this work. Third, we view our ARC method as new form of composite likelihood using all pairs, all triples and indeed all subsets of observations within each of the rows, and also within each of the columns. It thus captures all the data dependencies already included in pairwise likelihood and more while only requiring three estimating equations and $\mathcal{O}(N)$ cost.
@@ 3 The All-Row-Column Method
3.1 Three misspecified probit models
Our ARC approach for scalable inference in high-dimensional probit models with crossed-random effects combines estimates for three misspecified probit models. Each of them is constructed through omission of some random effects.
Consider the reparameterization $\bm{\psi}=(\bm{\gamma}^{\mathsf{T}},\tau^{2}_{A},\tau^{2}_{B})^{\mathsf{T}},$ whose components were already anticipated in Section 1, namely,
$\bm{\gamma}=\frac{\bm{\beta}}{(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}},\quad\tau^{2}_{A}=\frac{\sigma^{2}_{A}}{1+\sigma^{2}_{B}}\quad\text{and}\quad\tau^{2}_{B}=\frac{\sigma^{2}_{B}}{1+\sigma^{2}_{A}}.$ (7)
The proposed method begins with estimation of $\bm{\gamma}$ from the naïve model (3) that omits both of the random effects through maximization of the likelihood of the probit model
$L_{\text{all}}(\bm{\gamma})=\prod_{(i,j)\in\mathcal{S}}\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})^{1-y_{ij}}.$ (8)
Given $\hat{\bm{\gamma}}$, the maximizer of (8), we proceed with estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$ from the two probit models (4) and (5) that each omit one of the random effects. Accordingly, parameter $\tau^{2}_{A}$ is estimated from the row-wise likelihood

### v4

# Consistent and Scalable Composite Likelihood Estimation of Probit Models with Crossed Random Effects

@@ Abstract
Estimation of crossed random effects models commonly requires computational costs that grow faster than linearly in the sample size $N$, often as fast as $\Omega(N^{3/2})$, making them unsuitable for large data sets. For non-Gaussian responses, integrating out the random effects to get a marginal likelihood brings significant challenges, especially for high dimensional integrals where the Laplace approximation might not be accurate. We develop a composite likelihood approach to probit models that replaces the crossed random effects model with some hierarchical models that require only one-dimensional integrals. We show how to consistently estimate the crossed effects model parameters from the hierarchical model fits.We find that the computation scales linearly in the sample size. We illustrate the method on about five million observations from Stitch Fix where the crossed effects formulation would require an integral of dimension larger than $700{,}000$.
Keywords: Adaptive Gauss-Hermite Quadrature; Binary Regression; E-Commerce; High-Dimensional Data
@@ 1 Introduction
In this paper we develop a new composite likelihood approach to handling probit models with two crossed random effects. Our initial motivation was to get point and interval parameter estimates at a computational cost that grows only linearly in the sample size, $N$. Standard algorithms for crossed random effects typically have superlinear cost, commonly $\Omega(N^{3/2})$, making them unsuitable for modern large data sets. A second issue is that the marginal likelihood in a crossed random effects model is an integral over $\mathbb{R}^{D}$ where $D$ is large enough to make the integration problem challenging. Our scalable estimation method replaces this $D$-dimensional integral by $D$ integrals of dimension one.
The common notation for mixed effects models combines fixed and random effects through a formula such as $X{\beta}+Z{b}$ using matrices $X$ and $Z$ of known predictors with unknown coefficient vectors ${\beta}$ and ${b}$ where ${b}$ happens to be random. This formulation is simple and elegant but it hides some extremely important practical differences. As discussed above, the crossed setting leads to one high-dimensional integral while the hierarchical one uses many low-dimensional integrals. For Gaussian responses we can use generalized least squares on the response vector, without explicitly solving an integral. Even there, the crossed setting is harder. A hierarchical model has a block diagonal covariance matrix for the response vector resulting in a linear cost. For unbalanced crossed random effects, generalized least squares typically has a superlinear cost.
@@
With the size $N$ of data sets growing rapidly it is not possible to use estimation methods with a cost of $\Omega(N^{3/2})$. Ideally the cost should be $O(N)$. We are motivated by electronic commerce problems with large data sets. A company might have customers $i=1,\dots,R$, to which it sells items $j=1,\dots,C$. It would then be interested in modeling how a response $Y_{ij}$ depends on some predictors ${x}_{ij}\in\mathbb{R}^{p}$. If they do not account for the fact that $Y_{ij}$ and $Y_{is}$ are correlated due to a common customer $i$ or that $Y_{ij}$ and $Y_{rj}$ are correlated due to a common item $j$, they will get an inefficient estimate. This flaw is less severe when $N$ is so large. What is very serious is that they will get unreliable standard errors for their estimates. In a setting where accounting for random effects is computationally impossible we expect that many users will simply ignore them, getting very naïve variance estimates and finding too many things significant.
@@
A typical feature of data in our motivating applications is very sparse and imbalanced sampling. Only $N\ll RC$ of the possible $({x}_{ij},Y_{ij})$ values are observed. There is generally no simple structure in the pattern of which $(i,j)$ pairs are observed. It is common for the data to have very unequal sampling frequencies in each of the row and column variables.
The scaling problem is easiest to describe for generalised least squares solutions to linear mixed models, based on results from Gao & Owen (2020). They note that the algorithm for generalised least squares involves solving a system of $R+C$ equations in $R+C$ unknowns which has cost $\Omega\{(R+C)^{3}\}$ in standard implementations. If $RC>N$ then $\max(R,C)>N^{1/2}$ and so $(R+C)^{3}>N^{3/2}$. The average number of observations per level is $N/(RC)$. This ratio is well below one in our motivating applications, and as long as it is $o(N^{1/3})$, the cost of standard algebra will be superlinear.
Standard Bayesian solutions run into a similar difficulty. For an intercept plus crossed random effects model, Gao & Owen (2017) show that the Gibbs sampler takes $\Omega(N^{1/2})$ iterations that each have a cost proportional to $N$, for $\Omega(N^{3/2})$ cost overall. Several other Bayesian approaches considered there also had difficulty.
There has been recent progress on scalable algorithms for crossed random effects problems improving both the frequentist and Bayesian approaches. For regression problems Ghosh et al. (2022a) replace standard equation solving by a backfitting algorithm that has $O(N)$ cost per iteration and they give conditions under which the number of iterations to convergence is $O(1)$ as $N\to\infty$. See also Ghandwani et al. (2023) for regression with random slopes. Papaspiliopoulos et al. (2020) use a collapsed Gibbs sampler and then give conditions under which it has linear cost in $N$ for the intercept-only crossed random effects regression model. Ghosh & Zhong (2021) do the same after weakening a stringent balance assumption.
We consider a binary response requiring a generalized linear mixed effects model with crossed random effects that incur the high dimensional integration problem mentioned above. There are fewer scalable solutions for this problem. There is a frequentist approach by Ghosh et al. (2022b) and a Bayesian approach by Papaspiliopoulos et al. (2023). The all-row-column probit given here is simpler than those two approaches and uses much weaker sampling assumptions.
Ghosh et al. (2022b) developed a penalized quasi-likelihood approach to logistic regression on fixed effects and two crossed random effects. That approach had iterations costing $O(N)$ each and empirically the number of iterations was $O(1)$. They used estimating equations from Breslow & Clayton (1993) which were based on work by Schall (1991) to maximize the marginal likelihood using a Laplace approximation. The quantity being estimated is not exactly the maximum likelihood estimate. It is a posterior mode corresponding to a not very informative prior, using some plugged-in weights, a quantity that goes back to Stiratelli et al. (1984). Penalized quasi-likelihood has a bias that can prevent it from being consistent. Even with just the intercept and one random effect, it requires the number of levels of that effect as well as the number of observations at each of those levels to diverge to infinity in order to get a consistent estimate. With sample sizes $R=N^{\rho}$ and $C=N^{\kappa}$, for $\rho,\kappa\in(0,1)$ they require $\max(\rho+2\kappa,2\rho+\kappa)<2$ and the observation probability for $(x_{ij},Y_{ij})$ can only vary over a narrow range.
Papaspiliopoulos et al. (2023) extend the collapsed Gibbs sampler to obtain scalable Bayesian inference in generalized linear mixed models with crossed random effects using a reparameterization called local centering. They include an intercept and $K\geq 2$ crossed random effects, while also discussing how fixed effects could be incorporated. Their approach requires a stringent balance assumption. For our data we would need $N_{i\text{\tiny$\bullet$}}=N/C$ for all rows $i=1,\dots,R$ and $N_{\text{\tiny$\bullet$}j}=N/R$ for all $j=1,\dots,C$. With this condition, their cost per iteration is $O(N+R+C)$ in our notation. The total cost is this cost per iteration times a relaxation time. For $K=2$ random effects and a discrete response like we have, they introduce an auxiliary relaxation time $T_{\text{aux}}$ and show in Theorem 1 that the cost is $O[\max(N,R+C)\min\{2N/(R+C),T_{\text{aux}}\}]$. Our problem has $\max(R,C)\ll N$ and then their cost is linear in $N$ if and only if $T_{\text{aux}}=O(1)$. They get $T_{\text{aux}}=O(1)$ when the observation pattern is uniformly distributed over all patterns with $N_{i\text{\tiny$\bullet$}}=N/C$ and $N_{\text{\tiny$\bullet$}j}=N/R$. Both $R$ and $C$ much grow linearly with $N$. They then require $N_{i\text{\tiny$\bullet$}}=O(1)$ and $N_{\text{\tiny$\bullet$}j}=O(1)$.
Our balance criteria are minimal. Our main results require $\max(\max_{i}N_{i\text{\tiny$\bullet$}},\max_{j}N_{\text{\tiny$\bullet$}j})/N=o(1)$. Both $R$ and $C$ can grow with $N$ at different rates. Our simulations, but not our theory, use a balance condition with nearly equal values among the $N_{i\text{\tiny$\bullet$}}$ and nearly equal values among the $N_{\text{\tiny$\bullet$}j}$.
We study the probit model because its Gaussian latent variable is a good match for Gaussian random effects. The probit and logit link functions are nearly proportional outside of tail regions (Agresti, 2002, pp. 246–247), and so they often give similar results.
@@
Crossed models also differ from hierarchical models in that it is only recently that the maximum likelihood estimate for generalized linear mixed models has been shown to be consistent for crossed random effects. This was accomplished by the subset argument of Jiang (2013) showing that the score equations have a root that is consistent for the parameter. One of the main open questions in generalized linear mixed models, like the ones we consider here, is at what rate do the estimators of the model parameters converge? Intuitively one might expect $O(N^{-1/2})$ or $O\{\min(R,C)^{-1/2}\}$, in accordance with the results valid for nested designs (Jiang et al., 2022). In some settings estimators for different parameters converge at different rates. See Jiang (2013) for some discussion and Lyu et al. (2024) for recent work assuming balanced sampling.
@@ 2 The All-Row-Column Method
2.1 Preliminaries
We consider two crossed random effects. There is a vector ${a}\in\mathbb{R}^{R}$ with elements $a_{i}$ and another vector ${b}\in\mathbb{R}^{C}$ with elements $b_{j}$. Conditionally on ${a}$ and ${b}$, the $Y_{ij}$ are independent with
$\mathrm{pr}(Y_{ij}=1\mid{a},{b})=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}+a_{i}+b_{j}),$ (1)
where $\Phi(\cdot)$ is the standard normal cumulative distribution function. In these models, random effects are typically assumed to be uncorrelated, ${a}\sim{N}(0,\sigma^{2}_{A}I_{R})$ independently of ${b}\sim{N}(0,\sigma^{2}_{B}I_{C}),$ where $I_{n}$ is the $n\times n$ identity matrix. See, for example, McCullagh & Nelder (1989, page 444). The probit model has a representation in terms of latent variables $\varepsilon_{ij}\sim{N}(0,1)$, via
$Y_{ij}=1\{{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}+a_{i}+b_{j}+\varepsilon_{ij}>0\},$ (2)
where $1\{E\}$ is the indicator function of the event $E$. The probability (1) and the likelihoods derived from it are all conditional on the values of ${x}_{ij}$.
We write $\mathcal{S}\subset\{1,\dots,R\}\times\{1,\dots,C\}$ for the set of $(i,j)$ pairs where $({x}_{ij},Y_{ij})$ was observed. We also work conditionally on $\mathcal{S}$. In our motivating applications the pattern of observation/missingness could be informative. Addressing that issue would necessarily require information from outside the data. Furthermore, the scaling problem is still the subject of current research even in the non-informative missingness setting. Therefore we consider estimation strategies without taking account of missingness.
The likelihood for ${\theta}=({\beta}^{\mathrm{\scriptscriptstyle T}},\sigma^{2}_{A},\sigma^{2}_{B})^{\mathrm{\scriptscriptstyle T}}$ is a cumbersome integral of size $R+C$,